# On Geometry and Topology of Hessian Manifolds

### v4

@@ Abstract
For a differentiable manifold $M$, a pair $(M,\nabla)$ is termed an affine manifold, if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. A Riemannian metric $g$ on $M$ is called a Hessian metric on $(M,\nabla)$, if $g$ can be expressed locally as the Hessian quadratic form $\nabla(df)$ of some smooth potential $f$. An affine manifold equipped with a Hessian metric is termed a Hessian manifold. We study the topology and geometry of Hessian manifolds.
Key topological constraints are derived, including the vanishing of the Euler characteristic for any compact Hessian manifold and infinite torsion-freeness of its fundamental group. We demonstrate that compact Hessian manifolds with an abelian fundamental group are homeomorphic to the flat torus. The theory of flat line bundles is developed to introduce the Picard group and the canonical class, which are used to prove that a curved compact orientable Hessian manifold must be a mapping torus. Curvature properties are then discussed. As an application, we prove that all compact Hessian-Einstein manifolds are flat.
A Hessian manifold $((M,\nabla),g)$ is said to be of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $g=\nabla\eta$. We place emphasis on Hessian manifolds of Koszul type, establishing an equivalence between their existence and that of positive flat line bundles, space-like Lagrangian embeddings, and hyperbolic affine structures. A duality with radiant manifolds, analogous to the classical Legendre transformation, is explored. We also develop a geometric uncertainty principle on Hessian manifolds of Koszul type, and classify all complete Hessian surfaces and closed Hessian $3$-folds.
keywords:
Hessian metric, Affine differential geometry, regular convex domain, Flat line bundle, Legendre transform, Lagrangian embedding, para-Kähler manifold.
1 Basics on Hessian Geometry
@@ 1.1 Introduction and Background
The study of Hessian manifolds originates with the foundational work of J.-L. Koszul in 1961 [1]. While their roots lie in affine differential geometry, in recent decades, these geometric structures have garnered significant attention and emerged as indispensable tools across a diverse spectrum of mathematical and scientific disciplines. Their influence spans from the theoretical underpinnings of information geometry and complex analysis to practical applications in optimization and mathematical statistics, and even touches upon the abstract realms of homological mirror symmetry. This resurgence of interest highlights the deep connections these manifolds forge between seemingly disparate areas, offering a unifying geometric language to explore sophisticated phenomena.
In affine differential geometry and information geometry, Hessian metrics underpin the dually flat spaces used in the study of statistical manifolds, where the Amari–Chentsov tensor codifies divergences and dual connections. Hessian metrics also arise naturally in convex optimization, where barrier functions induce complete Hessian metrics on convex domains. For more details, we refer to [2] and [3].
Also, regarding the applications to mirror symmetry, the seminal work of Hitchin [4], Gross–Siebert [5], and Kontsevich–Soibelman [6] has shown that affine manifolds with singularities and Hessian metrics model the base of torus fibrations in the Strominger-Yau-Zaslow conjecture.
@@
However, despite this rich tapestry, fundamental questions about Hessian manifolds remain open, particularly regarding the interplay between global topology, curvature properties, and algebraic obstructions. In this manuscript, we undertake a study of Hessian manifolds from both topological and geometric perspectives:
After briefly summarizing the history and background of the theory of Hessian manifolds, we recall the definitions of affine atlases and flat connections before formally defining Hessian metrics. While the literature contains several well-known examples of Hessian metrics from information theory, such as the canonical metrics on the probability simplexes [11] and the family of Gaussian distributions [11], this collection has remained quite limited and tied to specific contexts. To address this, we introduce a powerful and explicit warped product formula (Theorem 1.8), a systematic construction that generates a vast new class of Hessian manifolds and demonstrates that structures like the Robertson-Walker solutions from general relativity are part of this geometric framework, significantly expanding the known examples beyond their original domains.
Our investigation then begins by establishing topological constraints that provide a rigid global framework for any Hessian structure. Foundational results by Shima and Yagi [12] in 1997 established that complete Hessian manifolds are aspherical, as their universal coverings are convex domains. Building on this, and on partial results by Kostant–Sullivan [13] in 1975 and Fried [14] in 1982, our work confirms Chern’s longstanding 1955 conjecture in this general setting by proving that the Euler characteristic of any compact Hessian manifold must vanish (Proposition 2.17). We also provide the first proof that the fundamental group of a complete Hessian manifold is torsion-free (Proposition 2.19), and give two different proofs of the fact that the fundamental group of a compact Hessian manifold is infinite (Corollary 2.7 and Remark 2.21). These new algebraic results possess immediate classificatory power, allowing us to prove that any compact Hessian manifold with an abelian fundamental group is homeomorphic to a $n$-torus (Theorem 2.22).
Having established this broad landscape, we switch from these foundational constraints to a more refined analysis by developing the theory of flat line bundles as our primary analytical tool. We introduce this as a novel framework for studying Hessian geometry and define two crucial new objects within it: the canonical class, an intrinsic invariant of the affine structure, and the locally constant Picard group. The latter classifies all flat $\mathbb{R}$-line bundles over a differentiable manifold up to gauge equivalence. As a direct consequence, we also establish that any orientable compact Hessian manifold with vanishing first Betti number must be flat (Theorem 2.45).
We then continue to study the natural question that in which case the Hessian metric admits a globally defined potential function. Using sheaf cohomology and spectral sequences technique, we show that a Hessian manifold admits a single global potential function if its fundamental group is finite (Theorem 2.4). Moreover, it is readily seen that a compact Hessian manifold cannot support a global potential function. This problem brings a historically significant class of Hessian manifolds into sharp focus: those of Koszul type, which admit a global 1-form potential. For this class, we prove using Hodge theory that the first Betti number of a Hessian manifold of Koszul type is always positive (Proposition 2.9).
A central achievement of our work is to reveal that Hessian metrics of Koszul type are part of a much larger network of connections than previously understood. Whilst the work of H.Shima [15] established an important equivalence between the existence of a Koszul-type Hessian metric and that of a hyperbolic affine structure, we establish a new, powerful list of equivalences, demonstrating that these conditions are also entirely equivalent to two other seemingly disparate properties: the existence of a positive flat line bundle (Proposition 2.50) and admitting a space-like Lagrangian embedding (Corollary 3.32) into its own cotangent bundle, equipped with the Riemann extension metric first studied by Patterson and Walker in 1952 [16], thereby building a new bridge between affine differential geometry and para-Kähler geometry.
To get a finer understanding of the moduli of hyperbolic affine structures on an oriented Hessian manifold of Koszul type, we define canonical class as the de-Rham cohomology class corresponding to the canonical line bundle. In 1968, J.-L.Koszul proved that the hyperbolicity is stable under small perturbations of the affine structure [17]. Using group cohomology with Lie algebra coefficients, we show that any hyperbolic affine structure can be deformed into another of which canonical class is rational (Theorem 3.50).
This cohesive picture is further enriched by developing a duality with radiant affine manifolds, i.e. differentiable manifolds of which transition functions are linear, which builds upon the classical theory of dual connections in information geometry pioneered by Amari and Chentsov [2]. As a new contribution, we provide an explicit expression for the $1$-form potential of the dual of a radiant Hessian manifold of Koszul type, resembling the classical Legendre transformation in theoretical mechanics, and use this to situate the celebrated Cheng–Yau metric [18] within our framework.
Next, in a digression section, we also develop a new link to pseudoconvexity in complex analysis of several variables by showing that the Sasaki lifting of a complete Hessian manifold is a Stein manifold (Proposition 3.5), and correspondingly, positive flat line bundles are lifted to positive holomorphic line bundles (Lemma 3.2). A further observation then yields a stronger real analytic embedding theorem for this class of manifolds (Theorem 3.6) of which dimension bound is more optimal than the classical Whitney embedding.
The synthesis of these tools now culminates in powerful structural theorems. By applying D.Tischler’s classical 1970 theorem on fibrations [19] to the first Koszul form, we prove that any curved compact orientable Hessian manifold must be a mapping torus (Theorem 2.48), revealing a remarkably rigid underlying structure, particularly in low dimensions (Corollary 5.15). This result is further strengthened by a de Rham-type splitting theorem, which shows that the universal cover of a compact Hessian manifold isometrically splits as a product of another contractible Riemannian manifold and the real line (Proposition 4.8). Using this splitting structure and tools from geometric analysis, we are able to prove that a compact Hessian manifold cannot support a metric of negative curvature(Theorem 4.10). As yet another important application, we then prove that all compact Hessian-Einstein manifolds are flat (Theorem 4.13). An interesting geometric uncertainty principle is also established on Hessian manifolds, revealing links to quantum mechanics.
After obtaining a number of theorems, we recall that constructing explicit examples of non-trivial curved compact Hessian manifolds has been a significant challenge, with most known examples being flat Bieberbach manifolds. We therefore address this by explicitly constructing a non-trivial family of compact radiant Hessian manifolds of Koszul type that possess negative scalar curvature 1-1, suggesting that our theory is not vacuous. Furthermore, this careful investigation in low dimension finally results in the first complete topological classification of closed orientable Hessian 3-manifolds(Theorem 5.15), namely, they are always mapping tori of periodic mapping classes.
Finally, as one of the highlights of this article, we adopt our general theory to low dimensions. Applying classical algebraic topology techniques and using the solvability of Heisenberg group $H_{3}(\mathbb{R})$, we provide the first complete topological classification of all complete Hessian surfaces (Theorem 5.11). We also derive severe constraints on the intersection form for Hessian $4$-manifolds, and then prove that any Einstein metric on a closed Hessian $4$-manifold must be flat (Proposition 5.22). Last but not least, a novel characterization for potentials of $2$-dimensional flat metrics is established, showing that a pseudo-Riemannian metric
$g=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is flat if and only if the vector $(f_{xx},f_{xy},f_{yy})$ of second derivatives lies on a cone surface in the Euclidean space $\mathbb{R}^{3}$ (Proposition 5.2).
Before closing this section, we provide a concise road map of the present article.
This paper is organized as follows: Section 1 reviews fundamental definitions and introduces a warped product construction as first examples. Section 2 establishes key topological obstructions, such as the vanishing of the Euler characteristic and the infinitude of the fundamental group. We then develop the theory of flat line bundles, proving that curved compact orientable Hessian manifolds must be mapping tori. Section 3 explores the interaction with other geometric fields, establishing a central equivalence between Hessian metrics of Koszul type and several other structures, including positive line bundles and space-like Lagrangian embeddings. A duality via a generalized Legendre transform and a link to complex analysis are also discussed. Section 4 delves into the geometric analysis of Hessian metrics, studying their curvature tensors and proving that a compact Hessian manifold cannot support a metric of negative curvature, and that all compact Hessian-Einstein manifolds are flat. Finally, Section 5 applies these results to low dimensions, yielding a complete topological classification of complete Hessian surfaces and closed orientable Hessian 3-folds, and deriving strong constraints for Hessian 4-manifolds. An appendix on homotopy tori is included to support our classification of compact Hessian manifolds with Abelian fundamental groups, while another appendix is devoted to the geometric analysis techniques used in our study of curvature properties.
@@ 1.2 Affine and Hessian Manifolds
for all $\alpha\in I$, where $x_{\alpha}^{1},\ldots,x_{\alpha}^{n}$ are the local coordinates associated to the chart $\left(U_{\alpha},\varphi_{\alpha}\right)$. Moreover, the map $D\longmapsto\mathscr{A}_{D}$ is a bijection from the set of all flat and torsion-free connections on the tangent bundle $TM\rightarrow M$ to the collection of all maximal affine atlases of $M$, and the atlas $\mathscr{A}_{D}$ is termed the adapted affine atlas of $D$. Therefore, fixing an affine structure on $M$ is equivalent to specifying a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. In view of this fact, by a slightly abuse of notation we also refer to a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$ as an affine structure of $M$, and refer to a pair $(M,\nabla)$ as an affine manifold, if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$.
@@
Let $(M,\nabla)$ be an affine manifold. A Riemannian metric $h$ on $M$ is called a Hessian metric on $(M,\nabla)$, if for any point $p\in M$, there exists an open neighborhood $U$ of $p$ in $M$, such that $h|_{U}=\nabla(df)$ for some $f\in C^{\infty}(U)$. A pair $((M,\nabla),h)$ is termed a Hessian manifold, if $h$ is a Hessian metric on the affine manifold $(M,\nabla)$.
@@
All Pontryagin classes and rational Chern classes of an affine manifold must vanish.
@@
Chern’s conjecture is known to hold in various cases. For example, the 2-dimensional case is settled by John Milnor [20] in 1957. Moreover, in [21], Bruno Klingler proved that if a compact affine manifold admits a parallel nowhere vanishing top differential form then its Euler characteristic is zero. In latter sections we will also confirm this conjecture for all compact Hessian manifolds.
1.3 Warped Product Construction
In this section, we shall show that any warped product over the real line $\mathbb{R}$ with Hessian fibers is again Hessian, which allow us to construct new Hessian manifolds from known ones. In particular, we obtain that hyperbolic spaces and sufficiently small domains in round $n$-spheres are Hessian.
The following trivial observation serves as a base case in the recursive constructions of Hessian metrics.
Example 1.7.
Let $(M,g)$ be an one dimensional Riemannian manifold, and let $\nabla$ be an arbitrary connection on the tangent bundle $TM\rightarrow M$. Then, since $\Omega^{2}(M)=0$, the connection $\nabla$ is automatically flat and torsion-free. Moreover, by the theory of ordinary differential equations, we have that $g$ is always a Hessian metric on $(M,\nabla)$.
Now we shall state the fundamental construction theorem of Hessian warped products.
@@ 1.3 Warped Product Construction
Let $\Omega$ be a domain in $\mathbb{R}^{n}$, and let $\varphi=\varphi(x_{1},\dots,x_{n})\in C^{\infty}(\Omega)$. Let $f=f(t)\in C^{\infty}(\mathbb{R})$ be a smooth function with nowhere vanishing first derivative, so that in particular
  1. 1.
    $f$ admits an inverse $F\colon f(\mathbb{R})\rightarrow\mathbb{R}$, i.e. the smooth function satisfying $f\circ F=id$;
  2. 2.
    $y_{1}:=\operatorname{exp}(f)x_{1},\,\dots\,,y_{n}:=\operatorname{exp}(f)x_{n},\,y_{n+1}\equiv y:=\operatorname{exp}(f)$ form a coordinate system of an open domain in the upper half space $\{(y_{1},\dots,y_{n},y)\in\mathbb{R}^{n+1}:y>0\}$.
Then, the smooth function
$\hat{\varphi}(y_{1},\dots,y_{n},y):=\varphi(y_{1}/y,\dots,y_{n}/y)y+\int_{0}^{y}F^{\prime}(\operatorname{log}(t))^{2}(y-t)t^{-2}dt$
is well-defined, and it holds that
$\sum_{i=1}^{n+1}\sum_{j=1}^{n+1}\frac{\partial^{2}\hat{\varphi}}{\partial y_{i}\partial y_{j}}dy_{i}\otimes dy_{j}=dt^{2}+\operatorname{exp}(f)\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}dx_{i}\otimes dx_{j}.$
@@
The proof of this theorem is a straightforward computation and hence omitted. ∎
Combining Example 1.7 and Theorem 1.8, one may conclude that many important metric tensors in geometry and theoretical physics are actually Hessian. Examples include but not limited to sufficiently small geodesic disks in space forms, and the Robertson–Walker solutions to Einstein’s field equation.
The following is just one particular example of many corollaries of the powerful theorem above.
Example 1.9.
On the upper half space $\mathbb{H}^{n}:=\{(x_{1},\dots,x_{n})\in\mathbb{R}^{n}:x_{n}>0\}$, consider the coordinates
$y_{1}:=\frac{2x_{1}}{x_{n}^{2}},\,\dots,\,y_{n-1}:=\frac{2x_{n-1}}{x_{n}^{2}},\,y_{n}:=\frac{1}{x_{n}^{2}}.$
Let $\nabla$ be the affine structure on $\mathbb{H}^{n}$ satisfying $\nabla(dy_{i})=0$ for all $i=1,\dots,n$. Define smooth function
$\varphi(y_{1},\dots,y_{n}):=\frac{y_{1}^{2}+\dots+y_{n-1}^{2}}{8y_{n}}-\frac{1}{4}\operatorname{log}(y_{n})$
on $\mathbb{H}^{n}$. Then the Hessian quadratic form of $f$ is precisely
$h:=\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\partial^{2}\varphi}{\partial y_{i}\partial y_{j}}dy_{i}\otimes dy_{j}=\frac{dx_{1}^{2}+\dots+dx_{n}^{2}}{x_{n}^{2}}.$
In particular, $((\mathbb{H}^{n},\nabla),h)$ is a Hessian manifold.
2 Global Invariants for Hessian manifolds
2.1 Obstructions from Sheaf Cohomology
In this section, we study the very natural question that in which case there exists a global potential $f\in C^{\infty}(M)$ such that $h=\nabla(df)$. Before we can give an answer to this question, we need to state and prove the following two lemmata from algebraic topology and sheaf theory.

### v6

# On Topology of Compact Hessian Manifolds

@@ Abstract
For a differentiable manifold $M$, a pair $(M,\nabla)$ is called an affine manifold if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. A Riemannian metric $g$ on $M$ is said to be a Hessian metric on $(M,\nabla)$ if the $(0,3)$-tensor $\nabla g$ is totally symmetric.
We investigate general topological properties of compact Hessian manifolds, in particular confirm a long-standing conjecture of S.S.Chern in this special case, and then apply these constraints to low dimensions, providing a thorough topological classification of complete Hessian surfaces and closed orientable Hessian manifolds of dimension less than $5$.
Keywords: Hessian Metric, Affine Manifold, Low-Dimensional Topology
1 Basics on Hessian Manifolds
@@ 1.1 Introduction and Background
The study of Hessian manifolds originates with the foundational work of J.-L. Koszul in 1961 [18]. While their roots lie in affine differential geometry, in recent decades, these geometric structures have attracted significant attention and emerged as indispensable tools across a diverse spectrum of mathematical and scientific disciplines. Their influence spans from the theoretical underpinnings of information geometry and complex analysis to practical applications in optimization and mathematical statistics, and even touches upon the abstract realms of homological mirror symmetry. This resurgence of interest highlights the deep connections these manifolds forge between seemingly disparate areas, offering a unifying geometric language to explore sophisticated phenomena.
In information geometry, the parameter space of a multi-dimensional smooth family of probability density functions is oftentimes a Hessian manifold, with its metric defined by the relative entropy, and the Amari–Chentsov tensor codifies its divergence function. Hessian metrics also arise naturally in convex optimization, where barrier functions induce complete Hessian metrics on convex domains. For more details, we refer to [3] and [20].
Also, for applications to mirror symmetry, the seminal work of Hitchin [14], Gross–Siebert [11], and Kontsevich–Soibelman [17] has shown that compact Hessian manifolds model the base of torus fibrations in the Strominger-Yau-Zaslow conjecture.
@@
However, despite this rich tapestry, fundamental questions about Hessian manifolds remain open, particularly regarding the interplay between global topology and algebraic obstructions. In this manuscript, we study Hessian manifolds from topological perspectives:
After briefly summarizing the history and background of the theory of Hessian manifolds, our investigation begins by answering the natural question that in which case a Hessian manifold $((M,\nabla),h)$ may admit a global potential $f\in C^{\infty}(M)$ such that $h=\nabla(df)$. Using techniques from sheaf cohomology and spectral sequences, we proved that A Hessian manifold admits a global potential if its fundamental group is finite (Theorem 2.4). However, foundational results by Shima and Yagi [24] in 1997 established that complete Hessian manifolds are aspherical, as their universal coverings are convex domains, which then immediately implies that the fundamental group of a compact Hessian manifold is infinite and torsion-free (Corollary 2.6 and Proposition 2.16). This contrast brings a historically significant class of Hessian manifolds into sharp focus: those of Koszul type, which admit global 1-form potentials.
Results from the literature [23] manifest that a compact affine manifold admits a Hessian metric of Koszul type if and only if its universal covering is developed into a salient convex cone. Compact affine manifolds of this kind are traditionally said to be hyperbolic. Utilizing tools from convex analysis, we prove a central structural result: any compact Hessian manifold with dimension no greater than $6$ is finitely covered by the product of a flat torus and a hyperbolic affine manifold (Theorem 4.15):
Theorem 1.
[…]
$((E,f^{*}\nabla),f^{*}h)=((B\times\mathbb{T}^{n-k},\hat{\nabla}\oplus D),g+ds^{2})$
[…]
This result is complemented by powerful fibration theorems, which state that any compact orientable Hessian manifold must be a mapping torus or a Bieberbach manifold (Proposition 4.2) and any compact hyperbolic affine manifold must admit a fibration over $\mathbb{S}^{1}$ with periodic monodromy (Corollary 4.12):
Theorem 2.
Let $(M,\nabla)$ be a compact oriented hyperbolic affine manifold. Then there exists a fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$ with connected and oriented fiber, such that the monodromy of the mapping torus $f\colon M\rightarrow\mathbb{S}^{1}$ is periodic.
Finally, as one of the highlights of this article, we apply these general constraints to low dimensions. We provide the first topological classification of all complete Hessian surfaces (Theorem 5.6) and topological classification of closed orientable Hessian 3-manifolds and 4-manifolds (Theorem 5.9 and Theorem 5.15):
Theorem 3.
Let $((M,\nabla),h)$ be a complete Hessian surface. Then $M$ is diffeomorphic to a complete flat Riemannian manifold of dimension two.
Theorem 4.
Let $((M,\nabla),h)$ be a closed orientable Hessian $3$-manifold. Then there exists a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ such that $M$ is diffeomorphic to the mapping torus of $\varphi$, where $\Sigma_{g}$ is a closed orientable surface of genus $g\geq 1$.
Theorem 5.
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. Then $M$ is diffeomorphic to one of the following smooth manifolds:
  1. 1.
    a Bieberbach 4-manifold;
  2. 2.
    the mapping torus of an isometry of a hyperbolic 3-manifold;
  3. 3.
    a $\Sigma_{g}$-bundle over $\mathbb{T}^{2}$ with periodic monodromy,
where $\Sigma_{g}$ here is the closed orientable surface of genus $g>1$.
Conversely, we also construct explicit examples, to show that indeed all $3$-manifolds mentioned above do admit Hessian structures (Proposition 5.7 and Proposition 5.8). Similar constructions in dimension four is left to the reader.
As an easy corollary of our fibration theorems, our work confirms Chern’s longstanding 1955 conjecture in this particular setting by proving that the Euler characteristic of any compact Hessian manifold must vanish (Proposition 4.3).
While our primary lens is topological, we also explore necessary analytical and geometric tools, including the Cheng-Yau metric, and a novel duality between Koszul type and radiant Hessian manifolds, resembling the classical Legendre transform in theoretical mechanics.
@@ 1.2 Affine and Hessian Manifolds
for all $\alpha\in I$, where $x_{\alpha}^{1},\ldots,x_{\alpha}^{n}$ are the local coordinates associated to the chart $\left(U_{\alpha},\varphi_{\alpha}\right)$. Moreover, the map $D]\longmapsto\mathscr{A}_{D}$ is a well-defined bijection from the set of all gauge equivalent classes of flat and torsion-free connections on the tangent bundle $TM\rightarrow M$ to the collection of all maximal affine atlases of $M$, and the atlas $\mathscr{A}_{D}$ is termed the adapted affine atlas of $D$. Therefore, fixing an affine structure on $M$ is equivalent to specifying a gauge equivalent class of flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. In view of this fact, by a slightly abuse of notation we also refer to a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$ as an affine structure of $M$, and refer to a pair $(M,\nabla)$ as an affine manifold, if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$.
@@
Let $(M,\nabla)$ be an affine manifold. A Riemannian metric $h$ on $M$ is called a Hessian metric on $(M,\nabla)$, if for any point $p\in M$, there exists an open neighborhood $U$ of $p$ in $M$, such that $h|_{U}=\nabla(df)$ for some $f\in C^{\infty}(U)$; or equivalently, the $(0,3)$-tensor $\nabla h$ is totally symmetric. A pair $((M,\nabla),h)$ is termed a Hessian manifold, if $h$ is a Hessian metric on the affine manifold $(M,\nabla)$.
@@
All Pontryagin classes of an affine manifold must vanish.
@@
Chern’s conjecture is known to hold in various cases. For example, the 2-dimensional case is settled by John Milnor [22] in 1957. Moreover, in [16], Bruno Klingler proved that if a compact affine manifold admits a parallel nowhere vanishing top differential form then its Euler characteristic is zero. In later sections we will also confirm this conjecture for all compact Hessian manifolds.
1.3 Concepts from Information Geometry
In this short section, we list several preliminary results that will become useful later.
We first recall the core concept of information geometry, that is, the notion of dual connections.
Definition 1.7.
$Zg(X,Y)=g\left(D_{Z}X,Y\right)+g\left(X,D_{Z}^{*}Y\right)$
holds for all smooth vector fields $X,Y,Z$ on $M$.
@@ 1.3 Concepts from Information Geometry
$\bar{\nabla}:=\frac{1}{2}\left(\nabla^{*}+\nabla\right)$
is the Levi-Civita connection of $(M,g)$.
@@
See [23] for a proof of this statement. ∎
Lemma 1.9.
Proof.
Take arbitrary vector fields $X,Y$ on $M$. Then
$(\nabla^{*}_{X}\theta)(Y)=X(\theta(Y))-\theta(\nabla^{*}_{X}Y).$
Now, since $\theta(Y)=g(\theta^{\sharp},Y)$, we have
Moreover, since
substituting this into the previous expression, we obtain
The following notion is also classical in information geometry.
Definition 1.10.
$A(X,Y,Z)=(\nabla_{Z}h)(X,Y)=h(\nabla^{*}_{X}Y-\nabla_{X}Y,Z)$
for all vector fields $X,Y,Z$ on $M$.
Corollary 1.11.
Let $((M,\nabla),h)$ be a compact orientable Hessian manifold, and let $A$ be its Amari–Chentsov tensor. Suppose that $(M,h)$ is a flat Riemannian manifold. Then it holds that $A\equiv 0$.
Proof.
Since $(M,h)$ is a flat Riemannian manifold, we have that $\nabla$ is identical to the Levi-Civita connection of $(M,h)$, and hence $A=\nabla h=0$ by metric compatibility. ∎
2 Topological Generalities
2.1 Methods of Sheaf Cohomology
In this section, we study the very natural question that in which case a Hessian manifold $((M,\nabla),h)$ may admit a global potential $f\in C^{\infty}(M)$ such that $h=\nabla(df)$. Before we can give an answer to this question, we need to state and prove the following two lemmata from algebraic topology and sheaf theory.